A meteorological balloon is inflated with helium gas to a volume of 4.5L at a pressure of 101.3kPa and a temperature of 300K (on earth). The balloon is released into the atmosphere. Determine the temperature at the new altitude if the new volume is 8.7L at a pressure of 35kPa.
1
Expert's answer
2020-05-19T08:47:34-0400

P1*V1/T1 = P2*V2/T2

T2 = T1*P2*V2/P1*V1 = 300K*35kPa*8.7L/(101.3kPa*4.5L) = 200.4 K